# Projected Augmented Waves (PAW): extended resolution of unity method

### v3

@@ Abstract
The Projected Augmented Waves (PAW) method is based on a linear transformation between the pseudo wavefunctions and the all electron wavefunctions. To obtain high accuracy with this method, it is important that the local part of the linear transform (inside each atomic sphere) be defined over a complete basis set (with deviations from completeness leading to corrections to the total energy not computed within current implementations of PAW). Here we show how to make this basis much closer to complete without significant additional computational work and without modifying the transformation in any significant way. This is done by extending the resolution of unity used for the transform to include more smooth wavefunctions (which have nothing to do with the atomic problem) and having them linearly transform via the identity.
@@ I Introduction
  • The calculation of eigenvalues and eigenvectors of the Khon Sham (KS) Hamiltonian often requires FLAPW (Full Potential Linearized Augmented Plane Waves [22, 16, 18, 21, 15, 19, 20, 17]) which requires a relatively complex setup with multiple steps and calculations. On the other hand full potential plane wave methods have simple expressions for KS Hamiltonian matrix elements and reduce to the identity matrix for wavefunction overlaps in many cases - in particular for norm conserving pseudopotentials.
@@
Norm conserving pseudopotentials are based on the idea that a perfectly spherical scattering potential (the atomic sphere, when small enough, is roughly spherically symmetric) can be represented by a scattering phase shift - in each partial wave and at each energy - as a boundary condition at the surfaces of the sphere. Any potential producing these scattering phase shifts is sufficient for calculations of eigenvalues of the KS Hamiltonian - not just the original atomic one. Therefore an equivalent problem where a much smoother potential with the same scattering phase shift (in the middle of the valence band) as the atomic potential, in each different angular momentum channel, can be used in the solution of the KS Hamiltonian. These smoother potentials may be used to solve for the eigenvalues of the KS problem leading to lower cutoffs (smoother wavefunctions) needed for electronic structure calculations. The norm conserving condition insures both accurate Madelung energy (with additional compensation charges for higher multipoles) and that the derivative of the scattering phase shift with respect to the energy is the same both for the atomic potential and the pseudopotential leading to improved accuracy as a function of energy (better transferability).
@@
A competitor to norm conserving pseudopotentials are ultrasoft potentials. These pseudopotentials use the many state generalizations of the Kleinman-Bylander transform [2] in order to reproduce the scattering phase shifts of the atomic problem simultaneously at many energies. This allows one to get rid of the norm conserving condition (needed for accuracy and transferability without multiple scattering energies) and restore correct Madelung energies and with a fictitious overlap matrix between the pseudized wavefunctions and compensating charges (for higher multipoles). This greatly softens the potential as the pseudopotential need not conserve norm and therefore has fewer constraints leading to softer pseudopotentials, especially for $3d$ transition metals which are hard to soften while conserving norm [25, 2, 38]. This also solves the problem of ghost bands: semi-core bands with significantly different scattering phase shifts then in the middle of the valence band that often plagued norm conserving pseudopotentials can now be handled by introducing a Kleinman-Bylander transform with a state at the energy of the semi-core states thereby insuring good wavefunctions there. The price for this is minor and associated with some additional terms in the KS equations which come about because of the fictitious overlap (norm).
@@
For PAW we introduce a linear transformation between the soft (low cutoff) pseudized wavefunctions and the all electron wavefunctions. The various quantities needed to set up the KS equation and obtain the ground state energy are then expressed through this linear transformation. PAW is sometimes considered the most versatile because at its core its an efficient method to map a pseudo-wave function problem onto an all electron problem - leading to great accuracy of computations. One of its limitations is that it assumes some completeness relations for pseudo wavefunctions used in the transform, a problem we here partially remedy.
@@ II PAW Summary
These relations are used to compute the form of the density of electrons the correlation and exchange and the Hartree (Coulomb piece) of the Hamiltonian [39, 2, 25]. This is essential to current PAW methodology [40]. Here we improve on this resolution of identity.
@@ III Main Construction
We now address the main problem pointed out in the previous section that the $\left|\tilde{\phi}_{i}\right\rangle$ do not form a complete basis set. This is in particular because the set $\left|\tilde{\phi}_{i}\right\rangle$ of the same cardinality as $\left|\phi_{i}\right\rangle$ and we must obtain $\left|\phi_{i}\right\rangle$ from atomic calculations, where typically only two or three wavefunctions, per angular momentum channel, are available with current PAW data sets [40]. Here we propose to circumvent this bottleneck by introducing:
@@
Here $\alpha$ runs over a list of on the order of ten additional pseudo wavefunctions per an all electron wavefunction. Where we have used the idea that it is much easier to generate smooth (low cutoff) wave functions - say by a combination of Bessel functions with different energies multiplied by spherical harmonics with different angular momentum channels, then to do atomic data set calculations. These wavefunctions can be chosen to have good overlaps with the plane waves (Bessel functions) in various angular momentum channels and therefore can be chosen as a nearly complete basis set for the space of the pseudized wave functions considered in the PAW method. As such multiple smooth wavefunctions can easily be added to the same all electron atomic calculation wavefunction. Where we have that:
@@ IV.2 Main modification to KS equations
The reason for the inequality is that we have used a different (more accurate) resolution of identity to simplify the situation which leads to different results, in particular note again the cross terms between $\alpha=1$ and $\alpha>1$. Similarly for
$\mathcal{T}_{M}^{\dagger}\hat{H}_{KS}\mathcal{T}_{M}=\hat{H}_{KS}+$
$\sum_{i,\alpha\in\mathbf{R}}\sum_{j\beta\in\mathbf{R}}\left|\tilde{p}_{i\alpha}\right\rangle\left[\left\langle\phi_{i\alpha}\mid\hat{H}_{KS}\mid\phi_{j\beta}\right\rangle-\left\langle\tilde{\phi}_{i\alpha}\mid\hat{H}_{KS}\mid\tilde{\phi}_{j\beta}\right\rangle\right]\left\langle\tilde{p}_{j\beta}\right|$
@@
Again the resolution of identity changed the final result. Furthermore the resolution of identity is used repeatedly for finding the ground state energy (see Section IV.3) and therefore its derivative with respect to the density $V_{KS}$ the KS potential.
@@ IV.3 Correlation and exchange and Hartree piece
Where we have repeatedly used Eq. (29) so Eq. (19) is much better then Eq. (9).
@@ V Conclusion
In this work we have introduced a new version (extended basis set) of the PAW transformation (given in Eqs. (12), (13), (14) and (15)) which allows for an extended resolution of identity as part of the map between the pseudized problem and the all electron problem. This greatly improves the accuracy with which mapping reproduces the key PAW equation for expectation values local operators (given in Eq. (10)) as well as the various applications of that identity to exchange and correlation energies, Coulomb energies and the KS secular equation. This improves the accuracy of PAW at very limited computational cost. In the future it would be of interest to include strong correlations into the calculations in the form of LDA+U or LDA+DMFT or LDA+GA [42, 43] with this new resolution of identity for the PAW method.

### v4

@@ Abstract
The Projected Augmented Waves (PAW) method is based on a linear transformation between the pseudo wavefunctions and the all electron wavefunctions. To obtain high accuracy with this method, it is important that the local part of the linear transform (inside each atomic sphere) be defined over a complete basis set (with deviations from completeness leading to corrections to the total energy not computed within current implementations of PAW). Here we show how to make this basis much closer to complete without significant additional computational work and without modifying the transformation in any significant way thereby making the modifications we propose easy to implement in current electronic structure codes for PAW.. This is done by extending the resolution of unity used for the transform to include more smooth wavefunctions (which have nothing to do with the atomic problem) and having them linearly transform via the identity.
@@ I Introduction
  • For all electron methods the calculation of eigenvalues and eigenvectors of the Khon Sham (KS) Hamiltonian often requires FLAPW (Full Potential Linearized Augmented Plane Waves [22, 16, 18, 21, 15, 19, 20, 17]) which requires a relatively complex setup with multiple steps and calculations. On the other hand full potential plane wave methods have simple expressions for KS Hamiltonian matrix elements and reduce to the identity matrix for wavefunction overlaps in many cases - in particular for norm conserving pseudopotentials.
@@
Norm conserving pseudopotentials are based on the idea that a perfectly spherical scattering potential (the atomic sphere, when small enough, is roughly spherically symmetric) can be represented by a scattering phase shift - in each partial wave and at each energy - as a boundary condition at the surfaces of the sphere. Any potential producing these scattering phase shifts is sufficient for calculations of eigenvalues of the KS Hamiltonian - not just the original atomic one. Therefore an equivalent problem where a much smoother potential with the same scattering phase shift (in the middle of the valence band) as the atomic potential, in each different angular momentum channel, can be used in the solution of the KS Hamiltonian. These smoother potentials may be used to solve for the eigenvalues of the KS problem leading to lower cutoffs (smoother wavefunctions) needed for electronic structure calculations. The norm conserving condition insures both accurate Madelung energy and that the derivative of the scattering phase shift with respect to the energy is the same both for the atomic potential and the pseudopotential leading to improved accuracy as a function of energy (better transferability).
@@
A competitor to norm conserving pseudopotentials are ultrasoft potentials. These pseudopotentials use the many state generalizations of the Kleinman-Bylander transform [2] in order to reproduce the scattering phase shifts of the atomic problem simultaneously at many energies. This allows one to get rid of the norm conserving condition (needed for accuracy and transferability without multiple scattering energies) and restore correct Madelung energies and with a fictitious overlap matrix between the pseudized wavefunctions and compensating charges (for higher multipoles). This greatly softens the potential as the pseudopotential need not conserve norm and therefore has fewer constraints leading to softer pseudopotentials, e.g. especially for $3d$ transition metals which are hard to soften while conserving norm as the $3d$ wavefunctions have no oscillations to smooth out with a pseudized wavefunction [25, 2, 38]. This also solves the problem of ghost bands: semi-core bands with significantly different scattering phase shifts then in the middle of the valence band that often plagued norm conserving pseudopotentials can now be handled by introducing a Kleinman-Bylander transform with a state at the energy of the semi-core states thereby insuring good wavefunctions there. The price for this is minor and associated with some additional terms in the KS equations which come about because of the fictitious overlap (norm).
@@
For PAW we introduce a linear transformation between the soft (low cutoff) pseudized wavefunctions and the all electron wavefunctions. The various quantities needed to set up the KS equation and obtain the ground state energy are then expressed through this linear transformation. PAW is sometimes considered the most versatile because at its core its an efficient method to map a pseudo-wave function problem onto an all electron problem - leading to great accuracy of computations. One of its limitations is that it assumes some completeness relations for pseudo wavefunctions used in the transform, a problem we here partially remedy at very low computational cost and in a manner that is highly compatible with current electronic structure codes for PAW.
@@ II PAW Summary
These relations are used to compute the form of the density of electrons the correlation and exchange and the Hartree (Coulomb piece) of the Hamiltonian [39, 2, 25]. This is essential to current PAW methodology [40]. Here we improve on this resolution of identity without changing its mathematical form or incurring significant additional computational costs.
@@ III Main Construction
We now address the main problem pointed out in the previous section that the $\left|\tilde{\phi}_{i}\right\rangle$ do not form a complete basis set. This is in particular because the set $\left|\tilde{\phi}_{i}\right\rangle$ of the same cardinality as $\left|\phi_{i}\right\rangle$ and we must obtain $\left|\phi_{i}\right\rangle$ from atomic calculations, where typically only two or three wavefunctions, per angular momentum channel, are available with current PAW atomic data sets [40]. Here we propose to circumvent this bottleneck by introducing:
@@
Here $\alpha$ runs over a list of on the order of ten to one hundred additional pseudo wavefunctions per an all electron wavefunction. Where we have used the idea that it is much easier to generate smooth (low cutoff) wave functions - say by a combination of Bessel functions with different energies multiplied by spherical harmonics with different angular momentum channels, then to do atomic data set calculations. These wavefunctions can be chosen to have good overlaps with the plane waves (Bessel functions) used for the basis for the KS problem for PAW in various angular momentum channels and therefore can be chosen as a nearly complete basis set for the space of the pseudized wave functions considered in the PAW method. As such multiple smooth wavefunctions can easily be added to the same all electron atomic calculation wavefunction. Where we have that:
@@ IV Some technical details
In this section we present the main technical consequences of our new transform as the apply to current implementations of PAW. However we note that since the mathematical structUre of our resolution of identity is identical to the one usually used for PAW these technical consequences would already be coded in electronic structure codes for PAW and require no further work to implement.
@@ IV.2 Main modification to KS equations
The reason for the inequality is that we have used a different (more accurate) resolution of identity to simplify the situation which leads to different results, in particular note again the cross terms between $\alpha=1$ and $\alpha>1$. Similarly for:
$\mathcal{T}_{M}^{\dagger}\hat{H}_{KS}\mathcal{T}_{M}=$
$=\sum_{i,\alpha\in\mathbf{R}}\sum_{j\beta\in\mathbf{R}}\left[\left\langle\phi_{i\alpha}\mid\hat{H}_{KS}\mid\phi_{j\beta}\right\rangle-\left\langle\tilde{\phi}_{i\alpha}\mid\hat{H}_{KS}\mid\tilde{\phi}_{j\beta}\right\rangle\right]\times$
×$\times\left|\tilde{p}_{i\alpha}\right\rangle\left\langle\tilde{p}_{j\beta}\right|+\hat{H}_{KS}$
@@
Again the resolution of identity changed the final result. Here $\hat{H}_{KS}=\nabla^{2}+V_{KS}$ where $V_{KS}$ is the KS potential. Furthermore the resolution of identity is used repeatedly for finding the ground state energy (see Section IV.3) and therefore also its derivative with respect to the electron density which is the value of the KS potential.
@@ IV.3 Correlation and exchange and Hartree piece
Where we have repeatedly used Eq. (29) which is dependent on the resolution of unity (see Section IV.1) so Eq. (19) is much better then Eq. (9) for total energy calculations.
@@ V Conclusion
In this work we have introduced a new version (extended basis set) of the PAW transformation (given in Eqs. (12), (13), (14) and (15)) which allows for an extended resolution of identity as part of the map between the pseudized problem and the all electron problem. This greatly improves the accuracy with which mapping reproduces the key PAW equation for expectation values local operators (given in Eq. (10)) as well as the various applications of that identity to exchange and correlation energies, Coulomb energies and the KS secular equation. This improves the accuracy of PAW at very limited computational cost and because of the mathematical form of our equations is nearly identical to that of regular PAW it can readily be incorporated into modern electronic structure implementations of PAW. In the future it would be of interest to include strong correlations into the calculations in the form of LDA+U or LDA+DMFT or LDA+GA [42, 43] with this new resolution of identity for the PAW method.